# Deep Sparse Latent Feature Models for Knowledge Graph Completion

### v1

@@ 5 Analysis
To showcase the interpretability of our model, derived from SBM, we visualize the latent structure learned from the WN18RR and FB15k-237 datasets in Figure 4. For demonstration purposes, we use stick-breaking prior settings of $\alpha_{\text{qry}}=100$, $\alpha_{\text{ans}}=50$, and a truncation level of $K=64$. The sparse latent feature matrix $F_{\text{ans}}$ shows how entities are grouped into communities, where larger absolute values suggest stronger confidence in whether a node belongs to a specific community. For WN18RR, as illustrated in Figure 4a, more pronounced clustering is visible, with the left and right columns showing larger absolute values, while the middle columns are more moderate. In contrast, the FB15k-237 matrix exhibits more evenly distributed values across its columns.
In addition, incorporating a text encoder allows for a more in-depth understanding of the latent structure learned from a KG. We select several communities and their most significant entities from the FB15k-237 dataset for enhanced visualization, with their textual descriptions provided in Table 4. This integration of text features enables more intuitive and concrete observations of the uncovered communities, validating both the effectiveness and interpretability of our approach.
Cluster : County
County Wexford: County Wexford is a county in Ireland…
Cluster : Music
PJ Harvey: Polly Jean Harvey MBE is an English musician…
Italo disco: Italo disco is a genre of music which originated in Italy…
Talent manager-GB: A talent manager, also known as band manager…
@@ 6 Related work
Knowledge Graph Completion. To address the task of KGC, initial research has concentrated on developing effective scoring mechanisms to evaluate the plausibility of triples embedded in low-dimensional spaces. A pioneering approach in this area is knowledge graph embedding (KGE) Bordes et al. (2013); Yang et al. (2014); Schlichtkrull et al. (2018); Sun et al. (2019); Balažević et al. (2019), also known as embedding-based methods. Notably, TransE Bordes et al. (2013) is a representative model that interprets a relationship $r$ as a translation from the head entity $h$ to the tail entity $t$. Tucker Balažević et al. (2019) employs Tucker Decomposition Tucker (1966) to compute a smaller core tensor and a set of three matrices, each representing entity and relation embeddings separately. Recently, text-based KGC methods have incorporated textual descriptions of entities and relations, thus encoding them into a more expressive semantic space. Specifically, NTN Socher et al. (2013) simplifies entity representation by averaging its word embeddings. SimKGC Wang et al. (2022) integrates a contrastive learning framework with three negative sampling strategies, significantly improving KGC performance. However, these prevalent KGC methods assume that the existence of a triple in a KG solely depends on the entities and relation involved, often overlooking the intricate interconnections among communities.
@@
KGC methods that leverage neighborhood information. Graph Neural Networks (GNNs), especially Message Passing Neural Networks (MPNNs), have become essential tools for node representation learning in graphs, where they assume that similar neighborhood structures yield closer node representations. Notable MPNN-based KGC methods like RGCN (Schlichtkrull et al., 2018), CompGCN (Vashishth et al., 2019), and KBGAT (Nathani et al., 2019) have demonstrated strong KGC performance but have since been found to inadequately leverage neighborhood information (Zhang et al., 2022; Li et al., 2023b). Furthermore, GNN-based approaches generally do not incorporate community-level information for KG completion. Meanwhile, there are few KGC methods that leverage clustering features, such as CTransR (Lin et al., 2015) and EL-Trans (Yang et al., 2023). However, these models often struggle with poor KGC performance and lack an end-to-end design, limiting their applicability to modern KGs.
@@
Stochastic Blockmodels have demonstrated success in uncovering various latent structures, thereby enhancing link prediction. The stochastic blockmodel (SBM) Holland et al. (1983) assigns each node to a specific community, with the interconnections between nodes influenced by their community memberships. The mixed membership stochastic blockmodel (MMSB) Airoldi et al. (2008) introduces a multinomial indicator vector for node-community assignments, allowing for mixed membership communities. However, MMSB restricts nodes to a single cluster at any given time. The overlapping stochastic blockmodel (OSBM) Latouche et al. (2011) overcomes this limitation by utilizing a multi-Bernoulli distribution, enabling nodes to belong to multiple communities simultaneously. The latent feature relational model (LFRM) Miller et al. (2009) is a specific instance of OSBM that applies the Indian Buffet Process (IBP) prior to the assignment matrix. Traditional SBMs, however, are constrained in expressiveness and scalability due to their reliance on MCMC Miller et al. (2009) or variational inference Zhu et al. (2016) for learning latent variables. Recently, DGLFRM Mehta et al. (2019) employs a sparse variational autoencoder (VAE) framework for inference in SBMs, thereby extending their applicability to larger graphs. Nevertheless, DGLFRM struggles to handle graphs with tens of thousands of nodes or more, a common scenario in modern KGs.
@@ 7 Conclusion
In this paper, we propose DSLFM-KGC, a framework developed to learn sparse latent structural features for enhancing knowledge graph completion (KGC). Specifically, we introduce a novel generative model for KGs, based on stochastic blockmodels (SBMs), which dynamically uncover latent communities and improve triple completion performance. Additionally, a deep sparse variational autoencoder enables scalable inference and greater expressiveness. Extensive experiments on three benchmark datasets verify the superior performance of DSLFM-KGC in completing missing triples while maintaining interpretability. Despite the improvements in KGC performance, there is still a significant challenge in optimizing training efficiency. Future research will focus on learning more expressive latent representations while reducing computational overhead.
Appendix A Mathematical proofs
@@ A.1 The negative ELBO
$=D_{\mathrm{KL}}\left[q_{\phi}(\mathcal{H})||p_{\theta}(\mathcal{H})\right]-\mathbb{E}_{q}\left[\log p_{\theta}(X_{\text{qry}},X_{\text{ans}},A|Z_{\text{qry}},Z_{\text{ans}},W_{\text{qry}},W_{\text{ans}})\right]$
$=D_{\mathrm{KL}}\left[q_{\phi}(\mathcal{H})||p_{\theta}(\mathcal{H})\right]-\mathbb{E}_{q}\left[\log p_{\theta}(X_{\text{qry}}|Z_{\text{qry}},W_{\text{qry}})\right]$
$~{}~{}-\mathbb{E}_{q}\left[\log p_{\theta}(X_{\text{qry}}|Z_{\text{qry}},W_{\text{qry}})\right]-\mathbb{E}_{q}\left[\log p_{\theta}(X_{\text{ans}}|Z_{\text{ans}},W_{\text{ans}})\right]$
$~{}~{}-\mathbb{E}_{q}\left[\log p_{\theta}(A|Z_{\text{qry}},Z_{\text{ans}},W_{\text{qry}},W_{\text{ans}})\right]$
@@
To compute the reconstruction terms, such as $\mathbb{E}_{q}\left[\log p_{\theta}(\mathbf{x}_{hr}|\mathbf{z}_{hr},\mathbf{w}_{hr})\right]$, we use the cosine similarity between the embedding $\mathbf{e}_{hr}$ (Equation 14) and the decoded representation $\mathbf{g}_{hr}$ (Equation 16):
$\mathbb{E}_{q}\left[\log p_{\theta}(\mathbf{x}_{hr}|\mathbf{z}_{hr},\mathbf{w}_{hr})\right]=\cos(\mathbf{e}_{hr},\mathbf{g}_{hr})$ (20)
Note that, $\mathbf{v}_{hr},\mathbf{v}_{t},\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr}$ and are $K$-dimensional vectors, while $\mathbf{g}_{hr},\mathbf{g}_{t}\in\mathbb{R}^{D}$. The time required to compute $\mathcal{L}_{\text{KL}}(B)$, $\mathcal{L}_{\text{Recon}}(B)$ and $\mathcal{L}_{\text{Comp}}(B)$ in Equation A.1 is $\mathcal{O}(|B|\cdot C_{\text{KL}})$, $\mathcal{O}(|B|\cdot C_{\text{Recon}})$ and $\mathcal{O}(|B|\cdot C_{\text{Comp}})$, with space complexities $\mathcal{O}(|B|\cdot K)$, $\mathcal{O}(|B|\cdot D)$ and $\mathcal{O}(|B|\cdot D)$, respectively. Here, $C_{\text{KL}}$, $C_{\text{Recon}}$ and $C_{\text{Comp}}$ denote the complexity for evaluating the KL divergence, reconstruction and triple completion terms for a single triple. Thus, the total time and space complexity for computing A.1 are $\mathcal{O}(|B|\cdot(C_{\text{KL}}+C_{\text{Recon}}+C_{\text{Comp}}))$ and $\mathcal{O}(|B|\cdot D+|B|\cdot K)$.
In practice, we apply two different weighting coefficients to the KL and reconstruction losses to balance the learning objectives and reduce the risk of posterior collapse (Higgins et al., 2017):
@@
Regarding the KL terms, we adhere to the method described by Kingma & Welling (2013) for computing the KL divergences for two normal variables: $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{w}_{hr})||p_{\theta}(\mathbf{w}_{hr})\right]$ and $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{w}_{t})||p_{\theta}(\mathbf{w}_{t})\right]$. In the following sections, we derive the computation of the KL divergence for two Beta distributions, i.e., $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{v}_{hr})||p_{\theta}(\mathbf{v}_{hr})\right]$ and $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{v}_{t})||p_{\theta}(\mathbf{v}_{t})\right]$, as well as for two Bernoulli distributions, i.e., $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{hr})||p_{\theta}(\mathbf{z}_{hr}|\mathbf{v}_{hr})\right]$ and $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{t})||p_{\theta}(\mathbf{z}_{t}|\mathbf{v}_{t})\right]$.
A.2 The KL divergence of Beta distributions
@@
A.3 The KL divergence of Concrete distributions
To enable differentiable optimization, we utilize the binary Concrete distribution (Maddison et al., 2016) to obtain a continuous relaxation of the Bernoulli distribution (Equation 12). However, the KL divergence of two Concrete distributions, $q(y)$ and $p(y)$, is intractable. We resort to approximation using the Monte Carlo (MC) expectations:
@@ A.3 The KL divergence of Concrete distributions
According to Maddison et al. (2016), the logarithm of the probability density function for the Concrete distribution is given by:
@@
where $q_{\phi}(v_{hr,j})$ is defined in Equation 11. Then $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{hr})||p_{\theta}(\mathbf{z}_{hr}|\mathbf{v}_{hr})\right]$ is estimated using Equation A.3. The computation of $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{hr})||p_{\theta}(\mathbf{z}_{hr}|\mathbf{v}_{hr})\right]$ is implemented similarly.
A.4 Reparameterization
In our model, the expectations over Beta, Bernoulli and Normal distributions is approximated using differentiable Monte Carlo (MC) estimate, as required by SGVB (Kingma & Welling, 2013). Furthermore, to draw samples from these distributions, a reparameterization trick is needed to ensure effective differentiation. To sample Normal variables $\mathbf{w}_{hr}$ and $\mathbf{w}_{t}$, we follow the standard approach used in vanilla VAE (Kingma & Welling, 2013). For reparameterization of Beta variables $\mathbf{v}_{hr}$ and $\mathbf{v}_{t}$, we adopt the implicit differentiation method (Figurnov et al., 2018).
To draw discrete Bernoulli variables during training, we utilize the Gumble-max relaxation (Maddison et al., 2014; Jang et al., 2016) to achieve a continuous approximation. Specifically, the distribution used to reparameterize $z_{hr,k}$ aligns with a binary special case of the Concrete distribution (Maddison et al., 2016):

### v2

@@ 5 Analysis
5.1 How does clustering enhance KGC performance?
To evaluate how clustering enhances KGC performance, we categorize testing triples into disjoint groups based on the geodesic distance (shortest path distance) between the head and tail entities. We then compared the Hit@k metrics for each group against those of SimKGC. The triple distributions for WN18RR and FB15k-237 are depicted in Figure 7. It is important to note that, in the Wikidata5M and UMLS test sets, no head entity within a triple is connected to any entity other than its corresponding tail entity. As demonstrated in Figure 4, DSLFM-KGC exhibits superior predictive accuracy on the WN18RR dataset, especially for long-range connections. For instance, DSLFM-KGC achieves a higher Hit@1 at a distance of 4 (64.9 compared to 58.3) and an improved Hit@10 at a distance of 5 (81.6 compared to 76.1). This observation suggests that the method effectively captures global information through community clustering, thereby enhancing KGC performance. Results for the FB15k-237 dataset are presented in Appendix D.2, where the improvement in long-range accuracy is less pronounced.
5.2 Qualitative analysis on Interpretability
To showcase the interpretability of our model, derived from SBM, we visualize the learned latent structure for the WN18RR and FB15k-237 datasets in Figure 11. Figure 11a reveals a more distinct clustering pattern in the WN18RR latent structure, characterized by larger absolute values in the left and right columns and more moderate values in the central columns. In contrast, the FB15k-237 matrix exhibits a more evenly distributed pattern of values across its columns.
Additionally, integrating a text encoder facilitates a more nuanced understanding of the latent structure learned from a KG. To illustrate this, we select several communities and their most significant entities from the FB15k-237 dataset for enhanced visualization, with their textual descriptions provided in Table 7. This integration of text features enables more intuitive and concrete interpretations of the uncovered communities, thereby validating both the effectiveness and interpretability of our approach.
5.3 Discussion
This section provides a comparative analysis of our method and LLM-based KGC approaches, focusing on both performance enhancement and interpretability. LLM-based methods such as CP-KGC zhang2024making and KICGPT wei2024kicgpt typically require extensive example demonstrations to guide LLMs toward effective KGC and mitigate hallucinations, as evidenced by the results in Table 1. This reliance on densely connected graph contexts also extends to their interpretability. Furthermore, LLM-based methods generally struggle to scale to large KGs comprising millions of entities due to their prohibitive computational costs. However, contemporary real-world KGs are often vast and sparsely connected (or even rarely connected) wang2021kepler , a domain in which our model demonstrates superior performance.
@@ 6 Related work
Knowledge Graph Completion. To address the task of KGC, initial research has concentrated on developing effective scoring mechanisms to evaluate the plausibility of triples embedded in low-dimensional spaces. A pioneering approach in this area is knowledge graph embedding (KGE) bordes2013translating ; yang2014embedding ; schlichtkrull2018modeling ; sun2018rotate ; balavzevic2019tucker , also known as embedding-based methods. Notably, TransE bordes2013translating is a representative model that interprets a relationship $r$ as a translation from the head entity $h$ to the tail entity $t$. Recently, text-based KGC methods have incorporated textual descriptions of entities and relations, thus encoding them into a more expressive semantic space. Specifically, NTN socher2013reasoning simplifies entity representation by averaging its word embeddings. SimKGC wang2022simkgc integrates a contrastive learning framework with three negative sampling strategies, significantly improving KGC performance. However, these prevalent KGC methods assume that the existence of a triple in a KG solely depends on the entities and relation involved, often overlooking the intricate interconnections among communities.
@@
KGC methods that leverage neighborhood information. Graph Neural Networks (GNNs), especially Message Passing Neural Networks (MPNNs), have become essential tools for node representation learning in graphs, where they assume that similar neighborhood structures yield closer node representations. Notable MPNN-based KGC methods like RGCN (schlichtkrull2018modeling, ), CompGCN (vashishth2019composition, ), and KBGAT (nathani2019learning, ) have demonstrated strong KGC performance but have since been found to inadequately leverage neighborhood information (zhang2022rethinking, ; li2023message, ). Furthermore, GNN-based approaches generally do not incorporate community-level information for KG completion. Meanwhile, there are few KGC methods that leverage clustering features, such as CTransR (lin2015learning, ) and EL-Trans (yang2023knowledge, ). However, these models often struggle with poor KGC performance and lack an end-to-end design, limiting their applicability to modern KGs.
@@
Stochastic Blockmodels have demonstrated success in uncovering various latent structures, thereby enhancing link prediction. The stochastic blockmodel (SBM) holland1983stochastic assigns each node to a specific community, with the interconnections between nodes influenced by their community memberships. The mixed membership stochastic blockmodel (MMSB) airoldi2008mixed introduces a multinomial indicator vector for node-community assignments, allowing for mixed membership communities. However, MMSB restricts nodes to a single cluster at any given time. The overlapping stochastic blockmodel (OSBM) latouche2011overlapping overcomes this limitation by utilizing a multi-Bernoulli distribution, enabling nodes to belong to multiple communities simultaneously. The latent feature relational model (LFRM) miller2009nonparametric is a specific instance of OSBM that applies the Indian Buffet Process (IBP) prior to the assignment matrix. Traditional SBMs, however, are constrained in expressiveness and scalability due to their reliance on MCMC miller2009nonparametric or variational inference zhu2016max for learning latent variables. Recently, DGLFRM mehta2019stochastic employs a sparse variational autoencoder (VAE) framework for inference in SBMs, thereby extending their applicability to larger graphs. Nevertheless, DGLFRM struggles to handle graphs with tens of thousands of nodes or more, a common scenario in modern KGs.
@@ 7 Conclusion
In this paper, we introduce DSLFM-KGC, a novel framework designed to learn sparse latent structural features for improved knowledge graph completion (KGC). Our approach leverages a probabilistic model for KGs inspired by stochastic blockmodels (SBMs), which dynamically discovers latent communities within the graph and subsequently enhances triple completion accuracy. To facilitate scalable inference and enhance model expressiveness, we incorporate a deep sparse variational autoencoder. Through extensive experimentation on three benchmark datasets, we demonstrate the superior performance of DSLFM-KGC in completing missing triples while maintaining interpretability. Despite these advancements, optimizing training efficiency remains a critical challenge. Future research will focus on developing methods to learn more expressive latent representations while simultaneously reducing computational overhead.
Appendix A Clustering Benefits for KGC: An Example
In the KG illustrated in Figure 5, entities are organized into overlapping communities. To address the query concerning the relationship between Michael Jordan and Gregg Popovich, "acquaintance" presents itself as a likely candidate, inferable from the observed interconnections between the "NBA Players" and "NBA Coaches" communities. This inference is not readily attainable through textual information alone.
Appendix B Mathematical proofs
@@ B.1 The negative ELBO
$=D_{\mathrm{KL}}\left[q_{\phi}(\mathcal{H})||p_{\theta}(\mathcal{H})\right]-\mathbb{E}_{q}\left[\log p_{\theta}(\mathcal{O}|Z_{\text{qry}},Z_{\text{ans}},W_{\text{qry}},W_{\text{ans}})\right]$
$=D_{\mathrm{KL}}[q_{\phi}(\mathcal{H}_{\text{qry}})||p_{\theta}(\mathcal{H}_{\text{qry}})]+D_{\mathrm{KL}}[q_{\phi}(\mathcal{H}_{\text{ans}})||p_{\theta}(\mathcal{H}_{\text{ans}})]-\mathbb{E}_{q}\left[\log p_{\theta}(X_{\text{qry}},X_{\text{ans}},A|Z_{\text{qry}},Z_{\text{ans}},W_{\text{qry}},W_{\text{ans}})\right]$
$=D_{\mathrm{KL}}[q_{\phi}(\mathcal{H}_{\text{qry}})||p_{\theta}(\mathcal{H}_{\text{qry}})]+D_{\mathrm{KL}}[q_{\phi}(\mathcal{H}_{\text{ans}})||p_{\theta}(\mathcal{H}_{\text{ans}})]-\mathbb{E}_{q}\left[\log p_{\theta}(X_{\text{qry}}|Z_{\text{qry}},W_{\text{qry}})\right]$
$~{}~{}-\mathbb{E}_{q}\left[\log p_{\theta}(X_{\text{ans}}|Z_{\text{ans}},W_{\text{ans}})\right]-\mathbb{E}_{q}\left[\log p_{\theta}(A|Z_{\text{qry}},Z_{\text{ans}},W_{\text{qry}},W_{\text{ans}})\right]$
@@
To compute the reconstruction terms, we use the cosine similarity between the embedding $\mathbf{e}_{hr}$ (Equation 15) and the decoded representation $\mathbf{g}_{hr}$ (Equation 17).
Note that, $\mathbf{v}_{hr},\mathbf{v}_{t},\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr}$ and are $K$-dimensional vectors, while $\mathbf{g}_{hr},\mathbf{g}_{t}\in\mathbb{R}^{D}$. The time required to compute $\mathcal{L}_{\text{KL}}(B)$, $\mathcal{L}_{\text{Recon}}(B)$ and $\mathcal{L}_{\text{Comp}}(B)$ in Equation B.1 is $\mathcal{O}(|B|\cdot C_{\text{KL}})$, $\mathcal{O}(|B|\cdot C_{\text{Recon}})$ and $\mathcal{O}(|B|\cdot C_{\text{Comp}})$, with space complexities $\mathcal{O}(|B|\cdot K)$, $\mathcal{O}(|B|\cdot D)$ and $\mathcal{O}(|B|\cdot D)$, respectively. Here, $C_{\text{KL}}$, $C_{\text{Recon}}$ and $C_{\text{Comp}}$ denote the complexity for evaluating the KL divergence, reconstruction and triple completion terms for a single triple. Thus, the total time and space complexity for computing B.1 are $\mathcal{O}(|B|\cdot(C_{\text{KL}}+C_{\text{Recon}}+C_{\text{Comp}}))$ and $\mathcal{O}(|B|\cdot D+|B|\cdot K)$.
In practice, we apply two different weighting coefficients to the KL and reconstruction losses to balance the learning objectives and reduce the risk of posterior collapse [20]:
@@
Regarding the KL terms, we adhere to the method described by [27] for computing the KL divergences for two normal variables: $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{w}_{hr})||p_{\theta}(\mathbf{w}_{hr})\right]$ and $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{w}_{t})||p_{\theta}(\mathbf{w}_{t})\right]$. In the following sections, we derive the computation of the KL divergence for two Beta distributions, i.e., $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{v}_{hr})||p_{\theta}(\mathbf{v}_{hr})\right]$ and $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{v}_{t})||p_{\theta}(\mathbf{v}_{t})\right]$, as well as for two Bernoulli distributions, i.e., $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{hr})||p_{\theta}(\mathbf{z}_{hr}|\mathbf{v}_{hr})\right]$ and $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{t})||p_{\theta}(\mathbf{z}_{t}|\mathbf{v}_{t})\right]$.
B.2 The KL divergence of Beta distributions
@@
B.3 The KL divergence of Concrete distributions
To enable differentiable optimization, we utilize the binary Concrete distribution [36] to obtain a continuous relaxation of the Bernoulli distribution (Equation 13). However, the KL divergence of two Concrete distributions, $q(y)$ and $p(y)$, is intractable. We resort to approximation using the Monte Carlo (MC) expectations:
@@ B.3 The KL divergence of Concrete distributions
According to [36], the logarithm of the probability density function for the Concrete distribution is given by:
@@
where $q_{\phi}(v_{hr,j})$ is defined in Equation 12. Then $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{hr})||p_{\theta}(\mathbf{z}_{hr}|\mathbf{v}_{hr})\right]$ is estimated using Equation B.3. The computation of $D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{hr})||p_{\theta}(\mathbf{z}_{hr}|\mathbf{v}_{hr})\right]$ is implemented similarly.
B.4 Reparameterization
In our model, the expectations over Beta, Bernoulli and Normal distributions is approximated using differentiable Monte Carlo (MC) estimate, as required by SGVB [27]. Furthermore, to draw samples from these distributions, a reparameterization trick is needed to ensure effective differentiation. To sample Normal variables $\mathbf{w}_{hr}$ and $\mathbf{w}_{t}$, we follow the standard approach used in vanilla VAE [27]. For reparameterization of Beta variables $\mathbf{v}_{hr}$ and $\mathbf{v}_{t}$, we adopt the implicit differentiation method [16].
To draw discrete Bernoulli variables during training, we utilize the Gumbel-max relaxation [37, 23] to achieve a continuous approximation. Specifically, the distribution used to reparameterize $z_{hr,k}$ aligns with a binary special case of the Concrete distribution [36]: